### v1

@@ 3.2 Learning the value function
In this approach, we approximate $\omega(t,x)=e^{-L_{\eta^{*}}(t,x)}$, where $\eta^{*}$ is obtained via Algorithm 1, using a neural network approximator. From the proof of Theorem 2.2, we have that with the optimal Lagrange parameters
@@
We then exploit elicitability of conditional expectations to find the best neural network approximator for $\omega$ denoted $\omega[{\mathfrak{o}}]$ where ${\mathfrak{o}}$ are the parameters of the NN. Leveraging elicitability for estimating conditional functionals using deep learning has been used in [11, 8, 16] to avoid time consuming nested simulations. We refer the interested reader to [12] for elicitability in a statistical context and Appendix C in [16] for a discussion on conditional elicitability and its application in deep learning. Specifically, conditional elicitability implies that minimising the loss function
@@
$\omega(t,x)=\operatorname*{arg\,min}_{h}\;{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\int_{0}^{T}\left(h(t,X_{t})-e^{-\int_{t}^{T}(\varsigma(u,X_{u})+\eta_{0}^{*}\,g_{u})\,{\rm d}u-\eta_{1}^{*}\,f_{T}}\right)^{2}\;{\rm d}t\right]\,,$
where the minimum is taken over all square integrable functions $h\colon{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}_{+}$, for more details see [12] and [16] Appendix C.
In implementation, we use the empirical estimator of the expectation by generating simulations of $X$ under ${\mathbb{Q}}[{\bar{\mu}}]$, using the same methodology as in Section 3.1, and then approximate the Riemman integrals in (18) using the same time discretisation as in the simulation of $X$. The detailed steps for estimating $\omega$ using the elicitability methodology are provided in Algorithm 3.
@@
4       generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}}\dots\overline{W}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(X_{t_{0}}^{{[m]}}\dots X_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
5      use samples to approximate $(\zeta_{i}^{[m]}:=\big{(}\int_{t_{i}}^{T}\big{(}\varsigma(u,X_{u}^{[m]})+\eta_{0}^{*}\,g(u,X_{u}^{[m]})\big{)}\,{\rm d}u\big{)}_{i\in\{0,\dots,N-1\},m\in{\mathfrak{M}}}$;
@@
${\mathfrak{L}}\big{[}\omega[{\mathfrak{o}}]\big{]}:=\frac{1}{M}\sum_{\begin{subarray}{c}m\in{\mathfrak{M}}\\ i=0,\ldots,N-1\end{subarray}}\left(\omega[{\mathfrak{o}}](t_{i},X_{t_{i}}^{[m]})-e^{-\zeta_{i}^{[m]}-\eta_{1}^{*}\,f(X_{t_{N}}^{[m]})}\right)^{2}\Delta t;$
@@ 4 Applications
This section is devoted to illustrations of the proposed methodology of combining experts opinions. Section 4.1 compares the two deep learning algorithms on simulated examples while Section 4.2 provides an application to implied volatility (IV) smiles. For the IV smiles application, we combine three different models of IV smiles, estimated from real data, and calculate the minimal weighted KL model that satisfies a constraint on the average skewness of the IV smiles.
@@ 4.1 Simulation examples
Using the methodology of learning the optimal drift explained in Section 3.1 and Algorithm 2, we ran experiments where we start the training using the same initial neural network, however, the number of steps for the Euler discretisation is increased from 10 to 100 to 1000 steps. For simplicity we consider a one-dimensional process with volatility that is constant $\sigma=1$ and where the drifts under the two experts’ models are
@@
[…]
@@
These choices of constrains mean that the processes at time $T=1$ should lie with 90% probability within the interval $(0.8,1.2)$, that is ${\mathbb{Q}}(X_{1}\in(0.8,1.2))=0.9$, and that the average time the process spends below a barrier is 0.2, specifically ${\mathbb{E}}^{\mathbb{Q}}[\int_{0}^{1}{\mathds{1}}_{\{X_{u}<t\}}{\rm d}u]\allowbreak=0.2$. In Table 1 we reports the value of the constraints under the two experts models, the average drift model, and the optimal model.
Table 1: Terminal constraints under various measures. ${\mathbb{E}}^{\circ}$ indicates the expectation under the different probability measures indicated in the columns.
@@
We observe that the value of the constraints of the average drift model is not the average of the value of the constraints of the experts, this is because ${\mathbb{Q}}[{\bar{\mu}}]$ average the drifts of both experts’ models.
The top two panels of Figure 1 show the corresponding losses during training as well as a scatter plot of the learnt RN derivative for the three different time discretisations. When the number of time steps is small, the discrete approximation of the stochastic integrals cannot capture the optimal measure, generated by $\frac{d{\mathbb{Q}}[\theta_{\eta^{*}}]}{d{\mathbb{Q}}[{\bar{\mu}}]}$, path-by-path. This explains why the scatter plot concentrates more and more along the diagonal as the number of time steps increases. Further, the bottom panel shows sample paths generated by the same Brownian motions for all four models: ${\mathbb{P}}^{(1)}$, ${\mathbb{P}}^{(2)}$, ${\mathbb{Q}}[{\bar{\mu}}]$, and ${\mathbb{Q}}[\theta_{\eta^{*}}]$. We observe that at terminal time, the sample paths are concentrated in the interval (0.8, 1.2) and that the sample paths are shifted upwards to enforce both constraints.
Next we examine how the approach where we learn the drift Algorithm 2 and the approach where we learn the value function compare Algorithm 3 to one another. For this, we consider the same underlying models as above, but with the three constraints
@@
Figure 2 shows the learnt model in the top panel (both methods produce indistinguishable results), as well as the evolution of the constraint as training progresses. After about 1,000 iterations, both approaches lead to constraints that are well within acceptable errors ($\lesssim 1\times 10^{-3}$) – see the first two panels in the bottom of Figure 2. The right most figure in the bottom panel shows the histogram of the difference between the measure change $\frac{{\rm d}{\mathbb{Q}}[\theta^{*}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ and $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ where $\theta^{*}$ is the drift obtained from either learning the drift (Algorithm 2) or learning the value function (Algorithm 3). They have estimated means of $1.1\times 10^{-7}\pm 10^{-3}$ and $6.1\times 10^{-8}\pm 1.1\times 10^{-3}$, respectively. Finally, we estimate the (constrained) KL divergence (2), using Monte Carlo simulations of $10,000$ sample paths, where the optimal $\theta$ stems from Algorithm 2 or Algorithm 3. For Algorithm 2 the estimated (constrained) KL divergence is $1.683\pm 0.005$ and for learning the value function it is $1.660\pm 0.005$. Thus, as expected, learning the value function through elicitability produces an optimal drift that results in a lower KL divergence when compared with learning the drift directly.
Figure 2: (top) sample paths under the expert models, the average drift model, and the optimal measure. (bottom) The left and middle panels show the evolution of the expectations as the two learning algorithms proceed. The right panel shows the histogram of the difference between the Radon-Nikodym derivative of the learnt model minus $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$.
@@ 4.2 Implied volatility smiles
In this section we investigate an application to modelling the evolution of implied volatility (IV) smiles. For this we estimate three different models (the expert models) using three distinct time frames and obtain the optimal model with a constraint on the at-the-money skewness. We use daily data for fixed time-to-maturity (TTM) IV smiles from the WRDS database for the assets AMZN at TTM of 60 days over the period July 8, 2010 to December 31, 2021. The data consists of pairs of Delta ($\Delta$) – rather than strike – and IV ($\sigma^{IV}$), that is $(\Delta_{i},\,\sigma^{IV}_{t,i})_{i=1,\dots,17=:I}$, for days $t=1,\ldots,T=2,893$, where $\Delta_{i}=0.1,0.15,\dots,0.9$. We approximate the real-world dynamics of the IV smile using the approach in [7], where the authors consider the entire IV surfaces. Here, we restrict to IV smiles for simplicity. In a first stage, for each time $t$, we project the data onto a functional basis of normalised Legendre polynomials $\{L_{j}(\Delta)\}_{j=0,..,4=:d-1}$ of up to order $4$ using linear regression. That is
where $x_{t}:=(x_{t,1},\ldots x_{t,d})$. This results in a time-series of coefficients ($(\{x_{t,j}\}_{j=1,\dots,d})_{t=1,\dots,T}$, which are shown in Figure 3.
@@
with $\mu:{\mathds{R}}^{d}\to{\mathds{R}}^{d}$ and $\sigma:{\mathds{R}}^{d}\to{\rm{S}}^{d}_{++}$, by maximising the likelihood of the observed data subject to a probability integral transform (PIT) penalty – [7] demonstrates that including the PIT penalty significantly decreases model misspecification. By telescoping the likelihood, we have the following loss function
$-\sum_{t=1}^{T}\Bigg{[}\frac{1}{2\Delta t}\left(\mu(x_{t})\Delta t-\Delta x_{t}\right)^{\intercal}\Sigma^{-1}(x_{t})\left(\mu(x_{t})-\Delta x_{t}\right)+\frac{1}{2}\log\det\big{(}\Sigma(x_{t})\big{)}\Bigg{]}+\zeta\,\sum_{j=1}^{d}PIT_{j},$
@@
To train the neural-SDE we employ two feed forward neural networks with five hidden layers and SiLU activation in all but the output layer for $\mu(x)$ and $\Sigma(x)$ separately. For the output layer, we use a $10\tanh(\cdot)$ activation function to bound the outputs to the range $[-10,10]$ – this improves the speed of convergence of the loss minimisation and also ensures that the resulting dynamics have strong solutions to the associated SDE. For $\Sigma(x)$, we reshape the output of the neural network into a $5$-dimensional lower triangular matrix $U(x)$ and set $\Sigma(x)=U(x)\,(U(x))^{\intercal}+10^{-3}I$, where $I$ is the $5\times 5$ identity matrix, to ensure strict positive definiteness.
To obtain the three expert models, we retrain (19) with a neural SDE model using data restricted to one of $[0,\frac{1}{3}T]$, $(\frac{1}{3}T,\frac{2}{3}T]$, and $(\frac{2}{3}T,T]$, but only retrain the drifts $\mu(x)$. Note that the experts’ models are absolutely continuous with respect to each other, thus they need to share the same volatility $\Sigma(x)$. This results in three distinct (experts) models, whose drift are $\mu^{(1)}(x)$, $\mu^{(2)}(x)$, and $\mu^{(3)}(x)$, and have volatility $\Sigma(x)$.
These three expert models together with the weights $\pi=(0.2,0.2,0.6)$ are then used to approximate the barycentre model subject to the constraint ${\mathbb{E}}^{{\mathbb{Q}}}[\partial_{\Delta}\sigma_{\tau}^{IV}(\Delta)|_{\Delta=\frac{1}{2}}]=0.05$ using the elicitability approach, see Algorithm 3, over a time horizon of $[0,\tau]$ years (with $\tau=\frac{1}{12}$). This constraint imposes that on average, the at-the-money skewness (i.e., at $\Delta=\frac{1}{2})$ of the IV smile at terminal time is equal to $0.05$. Figure 4 displays the sample paths of the 5-dimensional coefficients of the Legendre polynomials under the three experts, the average drift model, and the optimal model. Using simulations of the coefficients under the different models, we then generate under each model sample paths of IV smiles for the time horizon $[0,\tau]$.
@@
Figure 5 show a sample path of IV smiles from each expert, the model ${\mathbb{Q}}[{\bar{\mu}}]$, and the learnt constrained barycentre model ${\mathbb{Q}}[\theta_{\eta^{*}}]$, for four different times $t\in[0,28]$ days. The plots also display the $10\%$ to $90\%$ band of for $10,000$ sample paths from the ${\mathbb{Q}}[\theta_{\eta^{*}}]$ model. We observe that for later days, the bands increase and the models deviate from one another. Moreover, the displayed sample IV smile under ${\mathbb{Q}}[\theta_{\eta^{*}}]$ becomes steeper at $\Delta=\frac{1}{2}$ to meet the constraint.

### v2

@@ 5.2 Learning the value function
In this approach, we approximate $\omega(t,x)=e^{-L_{\eta^{*}}(t,x)}$, where $\eta^{*}$ is obtained via Algorithm 1, using a neural network approximator. From Theorem 4.3, we have that with the optimal Lagrange parameters
@@
We then exploit elicitability of conditional expectations to find the best neural network approximator for $\omega$ denoted $\omega[{\mathfrak{o}}]$ where ${\mathfrak{o}}$ are the parameters of the NN. Leveraging elicitability for estimating conditional functionals using deep learning has been used in [13, 10, 23] to avoid time consuming nested simulations. We refer the interested reader to [14] for elicitability in a statistical context and Appendix C in [23] for a discussion on conditional elicitability and its application in deep learning. For example, estimates of the conditional expectation ${\mathbb{E}}[Y|\sigma(Z)]$, where $Y$ and $Z$ are ${\mathcal{F}}$-measurable random variables with ${\mathbb{E}}[Y]<+\infty$, and $\sigma(Z)$ denotes the sigma algebra generated by $Z$, can be obtained by minimizing the loss function
${\mathfrak{L}}:={\mathbb{E}}\big{[}(h(Z)-Y)^{2}\big{]}$
over all functions $h$ s.t., $h(Y)$ is square-integrable. Denoting the optimiser as $h^{*}$, we would have $h^{*}(Z)={\mathbb{E}}[Y|\sigma(Z)]$. In our context, conditional elicitability implies that minimising the loss function
@@
$\omega=\operatorname*{arg\,min}_{h}\;{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[\int_{0}^{T}\left(h(t,X_{t})-e^{-\int_{t}^{T}(\varsigma(u,X_{u})+\eta_{0}^{*}\,g_{u})\,{\rm d}u-\eta_{1}^{*}\,f_{T}}\right)^{2}\;{\rm d}t\right]\,,$
where the minimum is taken over all square integrable functions $h\colon{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}_{+}$; for more details see [14] and [23] Appendix C.
With the approximation of $\omega$ at hand, we obtain an estimate of the optimal drift using (70). Note that $\nabla_{x}L_{\eta}(t,X_{t})$ can be obtained using backpropagation of $L_{\eta}(t,X_{t})$ with respect to the input parameters $x$.
In implementation, we use the empirical estimator of the expectation by generating simulations of $X$ under ${\mathbb{Q}}[{\bar{\mu}}]$, using the same methodology as in Section 5.1, and then approximate the Riemann integrals in (79) using the same time discretisation as in the simulation of $X$. The detailed steps for estimating $\omega$ using the elicitability methodology are provided in Algorithm 3.
@@
4       generate batch of paths $({\overline{W}}_{t_{0}}^{{[m]}}\dots\overline{W}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ and $(\tilde{X}_{t_{0}}^{{[m]}}\dots\tilde{X}_{t_{N}}^{{[m]}})_{m\in{\mathfrak{M}}}$ from ${\mathbb{Q}}[{\bar{\mu}}]$;
5      use samples to approximate $(\zeta_{i}^{[m]}:=\big{(}\int_{t_{i}}^{T}\big{(}\varsigma(u,\tilde{X}_{u}^{[m]})+\eta_{0}^{*}\,g(u,\tilde{X}_{u}^{[m]})\big{)}\,{\rm d}u\big{)}_{i\in\{0,\dots,N-1\},m\in{\mathfrak{M}}}$;
@@
${\mathfrak{L}}\big{[}\omega[{\mathfrak{o}}]\big{]}:=\frac{1}{M}\sum_{\begin{subarray}{c}m\in{\mathfrak{M}}\\ i=0,\ldots,N-1\end{subarray}}\left(\omega[{\mathfrak{o}}](t_{i},\tilde{X}_{t_{i}}^{[m]})-e^{-\zeta_{i}^{[m]}-\eta_{1}^{*}\,f(\tilde{X}_{t_{N}}^{[m]})}\right)^{2}\Delta t;$
@@ 6 Applications
This section is devoted to illustrations of the proposed methodology of combining experts opinions. Section 6.1 compares the two deep learning algorithms introduced in Section 5 on simulated examples while Section 6.2 provides an application to implied volatility (IV) smiles. For the IV smiles application, we combine three different models of IV smiles, estimated from real data, and calculate the minimal weighted KL model that satisfies a constraint on the average skewness of the IV smiles.
@@ 6.1 Simulation examples
Using the methodology of learning the optimal drift explained in Section 5.1 and Algorithm 2, we ran experiments where we start the training using the same initial neural network, however, the number of steps for the Euler discretisation is increased from 10 to 100 to 1000 steps [^2] Computation is carried out on an Intel Xeon CPU E5-2630 v4 2.20GHZ, 10 Core machine equipped with an NVIDIA TITAN RTX GPU (launched in 2018). With a batch–size of 1,024 and 1,000 time steps, our implementation runs 25 iterations per second. Newer machines should find much faster computation times.. For simplicity we consider a one-dimensional process with volatility that is constant $\sigma=1$ and where the drifts under the two experts’ models are
@@
As both expert models are OU processes with bounded coefficients, the results of section Section 3.2 apply.
The time horizon is $T=1$, the experts weights are $\pi_{1}=\pi_{2}=\frac{1}{2}$. Simulated paths under the two expert models are displayed in Figure 3. We observe that expert 1’s model has an upward drift while expert 2’s model is mean-reverting and cyclical. Figure 3 further displays simulated paths under the average drift model, which combines both the upward drift of expert 1 and the mean-reverting and cyclical pattern of expert 2.
@@
These choices of constraints mean that the processes at time $T=1$ should lie with 90% probability within the interval $(0.8,1.2)$, that is ${\mathbb{Q}}(X_{1}\in(0.8,1.2))=0.9$, and that the average time the process spends below a barrier is 0.2, specifically ${\mathbb{E}}^{\mathbb{Q}}[\int_{0}^{1}{\mathds{1}}_{\{X_{u}<t\}}{\rm d}u]\allowbreak=0.2$. Sample paths under the constraint barycentre model are given in the bottom right panel of Figure 3. In Table 1 we reports the value of the constraints under the two experts models, the average drift model, and the optimal model ${\mathbb{Q}}[\theta_{\eta^{*}}]$, i.e. the neural network approximation of the constrained barycentre. We observe that the mean under the optimal measure for both terminal and running constraints are essentially zero — in contrast to their expectations under either of the experts’ models or the mean drift measure.
Table 1: Constraints under various measures. ${\mathbb{E}}^{\circ}$ indicates the expectation under the different probability measures indicated in the columns.
@@
We observe that the value of the constraints of the average drift model is not the average of the value of the constraints of the experts, this is because ${\mathbb{Q}}[{\bar{\mu}}]$ averages the drifts of both experts’ models.
The top two panels of Figure 3 show the corresponding losses during training as well as a scatter plot of the learnt RN derivative $\frac{{\rm d}{\mathbb{Q}}^{*}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$, using the representation in (77), versus the targeted (true) RN derivative $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ in (72) for the three different time discretisations — using the same underlying Brownian sample paths. If the learnt drift and the simulation are exact, then this scatter plot should be a straight line with slope $1$. When the number of time steps is small, the discrete approximation of the stochastic integrals cannot capture the optimal measure, generated by $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$, path-by-path. This explains why the scatter plot, concentrates more and more along the diagonal (with slope $1$) as the number of time steps increases, and provides strong evidence of the algorithm converges to the correct solution. Further, the bottom panel shows sample paths generated by the same Brownian motions for all four models: ${\mathbb{P}}^{(1)}$, ${\mathbb{P}}^{(2)}$, ${\mathbb{Q}}[{\bar{\mu}}]$, and ${\mathbb{Q}}[\theta_{\eta^{*}}]$. We observe that under the optimal measures ${\mathbb{Q}}[\theta_{\eta^{*}}]$ at terminal time, the sample paths are concentrated in the interval (0.8, 1.2) and that the sample paths are shifted upwards to enforce both constraints.
Next we examine how Algorithm 2, the approach where we learn the drift, and Algorithm 3, the approach where we learn the value function, compare. For this, we consider the same underlying models as above, but with the three constraints
@@
Figure 4 shows sample paths of the learnt model in the top panel (both methods produce indistinguishable results) along the two experts models and the average drift model, as well as the evolution of the constraints as training progresses in the bottom panels. The bottom left panel shows the constraints for the “learning the value function” algorithm, that is Algorithm 3 while the bottom middle panel displays the constraints for “learning the drift” algorithm, that is Algorithm 2. After about 1,000 iterations, both approaches lead to constraints that are well within acceptable errors ($\lesssim 1\times 10^{-3}$) — see the first two panels in the bottom of Figure 4. The right most figure in the bottom panel shows the histogram of the difference between the measure change $\frac{{\rm d}{\mathbb{Q}}[\theta^{*}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ and $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ where $\theta^{*}$ is the drift obtained from either learning the drift (Algorithm 2) or learning the value function (Algorithm 3). They have estimated means of $1.1\times 10^{-7}\pm 10^{-3}$ and $6.1\times 10^{-8}\pm 1.1\times 10^{-3}$, respectively. Denoting by $\delta_{1}$ the difference between the RN derivative obtained by Algorithm 2 and the target RN derivative, i.e. $\delta_{1}:=\frac{{\rm d}{\mathbb{Q}}[\theta^{*}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}-\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$, and by $\delta_{2}$ the difference between the RN derivative obtained by Algorithm 3 and the target RN derivative, we perform a Welch’s T-test for $\delta_{1}$ and $\delta_{2}$. The Welch’s T-test (which tests for equality of mean in two distributions with differing variances) provides a T-statistic of $-3.26\times 10^{-5}$ and corresponding p-value of $0.999974$, indicating we cannot reject the null that they are indistinguishable. Furthermore, the one-sample T-test for each $\delta_{1}$ and $\delta_{2}$ individually, with a null hypothesis of a mean of zero, each has T-statistic of $-4.62\times 10^{-5}$ and $-2.88\times 10^{-6}$, respectively, and corresponding p-values of $0.99996$ and $0.999998$, respectively. As both of the estimated means are statistically indistinguishable from zero, and one another, this provides confidence that both algorithms converged to the true RN derivative. Finally, we estimate the (constrained) KL divergence (5), using Monte Carlo simulations of $10,000$ sample paths, where the optimal $\theta$ stems from Algorithm 2 or Algorithm 3. For Algorithm 2 the estimated (constrained) KL divergence is $1.683\pm 0.005$ and for learning the value function, Algorithm 3, it is $1.660\pm 0.005$. We performed a one-sided Welch’s T-Test with null hypothesis that KL divergence from learning the value function has a lower mean than that from learning the drift which resulted in a t-statistic of $-326.61$ and p-value that was (machine) indistinguishable from 0. Hence, we have extremely strong evidence that Algorithm 3 produces a smaller KL divergence than Algorithm 2. We anticipate that learning the value function Algorithm 3 is better than the brute-force approach of learning the drift Algorithm 2, as the former makes use of the underlying mathematical derivation of the optimal value function, and thus the formula for the optimal drift.
Figure 4: (top) sample paths under the expert models, the average drift model, and the optimal measure. (bottom) The left and middle panels show the evolution of the constraints as the two learning algorithms proceed. The right panel shows the histogram of the difference between the RN derivative of the learnt models minus the target RN derivative $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$.
@@ 6.2 Implied volatility smiles
In this section we investigate an application to modelling the evolution of implied volatility (IV) smiles — recall that IV smiles refer to the implied volatility for European options with a fixed time-to-maturity as a function of strike. For this, we estimate three different models (the expert models) using three distinct time frames and obtain the optimal model with a constraint on the at-the-money skewness. We use daily data for fixed time-to-maturity (TTM) IV smiles from the WRDS database for the assets AMZN at TTM of 60 days over the period July 8, 2010 to December 31, 2021. The data consists of pairs of Delta ($\Delta$) — rather than strike — and IV ($\sigma^{IV}$), that is $(\Delta_{i},\,\sigma^{IV}_{t,i})_{i=1,\dots,17=:I}$, for days $t=1,\ldots,T=2,893$, where $\Delta_{i}=0.1,0.15,\dots,0.9$. We approximate the real-world dynamics of the IV smile using the approach in [9], where the authors consider the entire IV surfaces. Here, we restrict to IV smiles for simplicity. In a first stage, for each time $t$, we project the data onto a functional basis of normalised Legendre polynomials $\{L_{j}(\Delta)\}_{j=0,\ldots,4=:d-1}$ of up to order $4$ using linear regression. That is
where $x_{t}:=(x_{t,1},\ldots x_{t,d})$. This results in a time-series of coefficients ($(\{x_{t,j}\}_{j=1,\dots,d})_{t=1,\dots,T}$, which are shown in Figure 5. That is our underlying stochastic process that the agent aims at combining has dimension five.
@@
with $\mu:{\mathds{R}}^{d}\to{\mathds{R}}^{d}$ and $\sigma:{\mathds{R}}^{d}\to{\rm{S}}^{d}_{++}$, by maximising the likelihood of the observed data subject to a probability integral transform (PIT) penalty — for further details on neural SDE estimation and why the PIT penalty decreases model misspecification, see [9]. By telescoping the likelihood, the corresponding loss function for (80) is
$-\sum_{t=1}^{T}\Bigg{[}\frac{1}{2\Delta t}\left(\mu(x_{t})\Delta t-\Delta x_{t}\right)^{\intercal}\Sigma^{-1}(x_{t})\left(\mu(x_{t})\Delta t-\Delta x_{t}\right)+\frac{1}{2}\log\det\big{(}\Sigma(x_{t})\big{)}\Bigg{]}+\zeta\,\sum_{j=1}^{d}PIT_{j},$
@@
To train the neural-SDE, to estimate (80), we employ two feed forward neural networks with five hidden layers and SiLU activation in all but the output layer for $\mu(x)$ and $\Sigma(x)$ separately. For the output layer, we use a $10\tanh(\cdot)$ activation function to bound the outputs to the range $[-10,10]$ — this improves the speed of convergence of the loss minimisation and also ensures that the resulting dynamics have strong solutions to the associated SDE. For $\Sigma(x)$, we reshape the output of the neural network into a $5$-dimensional lower triangular matrix $U(x)$ and set $\Sigma(x)=U(x)\,(U(x))^{\intercal}+10^{-3}I$, where $I$ is the $5\times 5$ identity matrix, to ensure strict positive definiteness. As both the drift and covariance matrices are outputs of feed-forward neural networks, with smooth, bounded activation functions (SiLU and tanh), they are guaranteed to be bounded and Lipschitz. Consequently, 2.1 and 2.2, and the conditions in Theorem 4.3 are satisfied, so that we may apply our optimal measure change results. Furthermore, while we do not impose any static arbitrage constraints on the generated implied volatility smiles, [9] finds that generated smiles (in their case entire surfaces) have a lower percentage of static arbitrage than the raw data itself.
To obtain the three expert models, we retrain (80) with a neural SDE model using data restricted to one of $[0,\frac{1}{3}T]$, $(\frac{1}{3}T,\frac{2}{3}T]$, and $(\frac{2}{3}T,T]$, but only retrain the drifts $\mu(x)$. Note that, in our problem formulation, the experts’ models are absolutely continuous with respect to each other, thus they need to share the same volatility $\Sigma(x)$. This results in three distinct (experts) models, whose drift are $\mu^{(1)}(x)$, $\mu^{(2)}(x)$, and $\mu^{(3)}(x)$, and they share the volatility $\Sigma(x)$.
These three expert models together with the weights $\pi=(0.2,0.2,0.6)$ are then used to approximate the barycentre model subject to the constraint ${\mathbb{E}}^{{\mathbb{Q}}}[\partial_{\Delta}\sigma_{\tau}^{IV}(\Delta)|_{\Delta=\frac{1}{2}}]=0.05$ using the elicitability approach, see Algorithm 3, over a time horizon of $[0,\tau]$ years (with $\tau=\frac{1}{12}$). This constraint imposes that on average, the at-the-money skewness (i.e., at $\Delta=\frac{1}{2})$ of the IV smile at terminal time is equal to $0.05$ — which is slightly higher than the average at-the-money skewness of each of the expert’s model. We further place a higher weight on the expert model that is trained with more recent data, and equal weights on the other two expert models. Figure 6 displays the sample paths of the 5-dimensional coefficients of the Legendre polynomials under the three experts’ models (estimated via the neural SDE), the average drift model, and the optimal model. Using the simulations of the coefficients of the Legendre polynomials under the different models, we then generate, under each model, sample paths of IV smiles for the time horizon $[0,\tau]$.
@@
Figure 7 shows a sample path of IV smiles from each expert, the model ${\mathbb{Q}}[{\bar{\mu}}]$, and the learnt constrained barycentre model ${\mathbb{Q}}[\theta_{\eta^{*}}]$, for four different times $t\in[0,28]$ days. The plots also display the $10\%$ to $90\%$ band of for $10,000$ sample paths from the ${\mathbb{Q}}[\theta_{\eta^{*}}]$ model. We observe that for later days, the bands increase and the models deviate from one another. Moreover, the displayed sample IV smile under ${\mathbb{Q}}[\theta_{\eta^{*}}]$ becomes steeper at $\Delta=\frac{1}{2}$ to meet the constraint.